Kutztown University Department of Mathematics

Colloquium

11:00 a.m.

February 26, 2008

Lytle Hall 228

Oh the Places You’ll Go:

From a MapQuest to a MATHquest

Dr. Jennifer Gorman

Gannon University

Abstract

When taking a trip we are often concerned with how much the trip will cost. One aspect of a trip is the tolls on a road, so in planning our route we can try to travel on the cheapest roads available to us. In this talk we will use Graph Theory to help us explore more formally the idea of travelling on the cheapest roads. More specifically we will investigate a variant of the Travelling Salesman Problem as well as the Hamiltonian Path Problem for some special classes of graphs. We will also observe how often in Mathematics we set out to prove one result and along the way we find ourselves with a new problem or ”quest” we need to conquer.

10:50 a.m. refreshments served

11:00 a.m. talk begins

Kutztown Univ. of Pennsylvania Dept. of Mathematics Dr. Paul Ache, Chair